# Mini-batch Estimation for Deep Cox Models: Statistical Foundations and Practical Guidance

### v2

@@ 1 Introduction
Cox proportional hazards regression (Cox, 1972) is one of the most commonly used approaches in survival analysis, where the outcome of interest is the time to a certain event. It assumes that the covariates have linear effects on the log-hazard function. With the development of deep learning, Cox deep neural networks (Cox-NN, or deep Cox model) have been proposed to capture the potential nonlinear relationship between covariates and survival outcomes to improve survival prediction accuracy (Faraggi & Simon, 1995; Katzman et al., 2018; Ching et al., 2018). Despite the success of the Cox model in survival analysis, it faces a significant optimization challenge when applied to large-scale data. In particular, the Cox model is typically trained by maximizing the partial likelihood (Cox, 1975). The maximum partial likelihood estimator (MPLE) is obtained through the gradient descent (GD) algorithm, which requires the entire dataset to compute the gradient. This approach is computationally demanding and memory-intensive, especially with large datasets. For example, in our motivating application, where images are used to predict survival outcomes (as detailed in Section 6), performing GD in Cox-NN with high-dimensional predictors and a large sample size is infeasible due to hardware memory constraints. Even traditional Cox regression faces optimization challenges for large-scale data. Tarkhan & Simon (2024) reported that the GD algorithm for Cox regression is prone to round-off errors when dealing with large sample sizes. Therefore, the scalability of the Cox model is substantially limited by the inefficiencies of GD optimization when applied to large-scale data.
@@
The stochastic gradient descent algorithm (SGD) is a scalable solution for optimization with large-scale data and has been widely used for training NN (Amari, 1993; Bottou, 2012). SGD alleviates the computational and memory burdens by using a randomly selected (small) subset of the dataset, known as a mini-batch, to compute the gradient and update parameters in each iteration. Moreover, in NN optimization, Xie et al. (2020) showed that, compared to GD, SGD favors flat minima, which often leads to better generalization. This explains why SGD is preferred for training an NN. However, SGD cannot directly optimize the partial likelihood of all samples through mini-batches because evaluating the partial likelihood of an event subject requires access to data for all subjects who survived longer (i.e., the at-risk set).
@@
Several attempts have been made to enable parameter updates for Cox models using mini-batch data. Kvamme et al. (2019) followed the idea of nested case-control Cox regression from Goldstein & Langholz (1992), proposing an approximation of the gradient of partial likelihood using the case-control pairs instead of using all at-risk samples. Sun et al. (2020) fitted a Cox-NN to establish a prediction model for an eye disease progression through SGD, where the iteration was based on the partial likelihood of a random subset of data. Tarkhan & Simon (2024) studied SGD for Cox regression in an online learning setting and demonstrated that the gradient of the expected mini-batch partial likelihood at the true parameter is zero. Despite the successful application of SGD to Cox models, the statistical foundations of the mini-batch maximum partial likelihood estimator (mb-MPLE), of which the SGD seeks to estimate, remain unexplored. Notably, the average mini-batch partial likelihood depends on the batch size and differs from the full partial likelihood, as highlighted in (2.8). Such a distinction in the objective function differentiates mb-MPLE from MPLE, requiring the development of new statistical properties for the mb-MPLE.
@@
  • Secondly, we provide practical guidance for training Cox-NN via SGD to search for the mb-MPLE. In practice, both the learning rate and batch size used in SGD are important hyperparameters, as they significantly influence the SGD dynamics. In optimizations where the objective function is independent of batch size, such as empirical risk minimization, the ratio of the learning rate to the batch size has been identified as a key factor in SGD dynamics during NN training (Goyal et al., 2017; Jastrzebski et al., 2017; Xie et al., 2020). This observation guides a hyperparameter tuning strategy, where the learning rate or the batch size is fixed while tuning the other. However, for Cox-NN, we note that the function targeted by SGD is batch-size dependent, so it remains unclear whether these results can be applied. To this end, we investigate how the local convexity of the objective changes with the batch size and extend the applicability of the tuning strategy to Cox-NN optimization. We provide both theoretical insights and numerical evidence that the ratio of the learning rate to the batch size remains a crucial factor in SGD dynamics during Cox-NN training. This offers insight into Cox-NN hyperparameter tuning.
  • Lastly, we develop the statistical properties of mb-MPLE for Cox regression with linear covariate effects. We show that mb-MPLE is $\sqrt{n}$-consistent and asymptotically normal, with asymptotic variance depending on the batch size. Furthermore, based on both theoretical analysis and numerical experiments on the impact of batch size on local convexity, we demonstrate that doubling the batch size improves the statistical efficiency of mb-MPLE. This phenomenon is not observed in other SGD optimizations, such as empirical risk minimization, where the optimizer’s statistical efficiency is independent of the SGD batch size. Moreover, for Cox regression, we study the numerical convergence of the SGD algorithm to the mb-MPLE in terms of iterations. We note that the objective function optimized by SGD is not global strong convex, and an additional projection step is necessary (namely, the projected SGD). We follow the non-asymptotic analysis of the projected SGD from Moulines & Bach (2011) to demonstrate that the algorithm approximates the mb-MPLE given sufficiently many iterations.
@@ 2.1 Cox Model and Deep Neural Network
where the matrices $W_{k}\in\mathbb{R}^{p_{k+1}}\times\mathbb{R}^{p_{k}}$ and vectors $v_{k}\in\mathbb{R}^{p_{k+1}}$ (for $k=0,\dots,K$) are the parameters of the NN. The pre-specified activation function $\sigma(\cdot)$ is a nonlinear transformation that operates component-wise on a vector, that is, $\sigma((x_{1},\dots,x_{p_{k}})^{T})=(\sigma(x_{1}),\dots,\sigma(x_{p_{k}}))^{T}$, for $k=1,\dots,K$. For NN in (2.2), $K$ denotes the depth of the network. The sequence $\mathbf{p}$ lists the width of each layer with the first element $p_{0}$ being the dimension of the input variable, $p_{1},\dots,p_{K}$ are the dimensions of the $K$ hidden layers, and $p_{K+1}$ is the dimension of the output layer. The matrix entries $(W_{k})_{i,j}$ are the weight linking the $j$th neuron in layer $k$ to the $i$th neuron in layer $k+1$, and the vector entries $(v_{k})_{i}$ represent a shift term associated with the $i$th neuron in layer $k+1$. In Cox-NN, the output dimension $p_{K+1}=1$ since the output $f_{\theta}(X)$ is a real value. We consider the commonly used activation function named Rectified Linear Units (ReLU) in Nair & Hinton (2010), i.e., $\sigma(x)=\max\{x,0\}$, which has been similarly considered in Schmidt-Hieber (2020) and Zhong et al. (2022).
@@
We refer to the minimizer of (2.3) as the MPLE. In the partial likelihood, sample $i$ who has experienced the event ($\Delta_{i}=1$) is compared to all samples that have not yet experienced the event up to time $T_{i}$, known as the at-risk set. Minimizing (2.3) is typically solved using gradient-descent-based algorithms, where the gradient $\nabla_{\theta}L^{(n)}_{Cox}(\theta)$ is calculated using all $n$ samples at each iteration (Therneau et al., 2015; Katzman et al., 2018; Zhong et al., 2022). The consistency and asymptotic normality of the MPLE have been well-studied by Andersen & Gill (1982) for Cox regression. For Cox-NN, the consistency of the MPLE and its convergence rate have been developed by Zhong et al. (2022).
However, when optimizing the Cox model through the SGD algorithm using a subset of samples, the algorithm does not directly minimize (2.3), which will be further explained in the next section.
@@ 2.2 Stochastic Gradient Descent for Cox Model
where $\gamma_{t}$ is a pre-scheduled learning rate for the $t$-th iteration. The gradient ${\nabla_{\theta}}L^{(s)}(D_{t}(s);\hat{\theta}_{t})$ is calculated from the mini-batch data $D_{t}(s)$ at $t$-th iterations. Suppose $D_{t}(s)$ is randomly selected from $D(n)$ with ${\nabla_{\theta}}L^{(s)}(D_{t}(s);\hat{\theta}_{t})$ uniformly bounded for all $t$, when $t\to\infty$, it has been shown that SGD (2.4) with an appropriate learning rate will find the minimizer of $\mathbb{E}_{D_{t}(s)}[L^{(s)}(D_{t}(s);\hat{\theta}_{t})|D(n)]$ if it is strongly convex, where the expectation is over the randomness of generating $D_{t}(s)$ from $D(n)$ (Moulines & Bach, 2011; Toulis & Airoldi, 2017). Throughout the paper, we assume $D_{t}(s)$ is randomly sampled from $D(n)$ at each iteration, so its distribution is independent of $t$. To simplify the notation, for any $\theta$, we omit the input data and drop the subscript $t$ and let $L^{(s)}(\theta):=L^{(s)}(D(s);\theta)$, $L^{(n)}(\theta):=L^{(n)}(D(n);\theta)$, and $\mathbb{E}[L^{(s)}(\theta)|D(n)]:=\mathbb{E}_{D(s)}[L^{(s)}(D(s);\theta)|D(n)]$.
@@
For the Cox model, the SGD implementation (Sun et al., 2020; Tarkhan & Simon, 2024) requires at least two samples to calculate the gradient, which restricts $s\geq 2$. The parameters are updated through
@@ 3.1 Statistical Properties of the mb-MPLE
Zhong et al. (2022) studied the asymptotic properties of the MPLE which minimizes (2.3). They showed that the MPLE achieves the minimax optimal convergence rate up to a polylogarithmic factor. While our work is inspired by the theoretical developments in Zhong et al. (2022), it has significant differences. First, the at-risk term $\frac{1}{n}\sum_{j=1}^{n}I(T_{j}\geq t)\exp\{f_{\theta}(X_{j})\}$ converges to a fixed function $\mathbb{E}[I(Y>t)\exp\{f_{\theta}(X)\}]$. However, in the average mini-batch log-partial likelihood, the at-risk term in $L^{(s)}_{Cox}(\theta)$ always consists of $s$ sub-samples. Thus, both the empirical loss $\mathbb{E}[L^{(s)}_{Cox}(\theta)\lvert D(n)]$ and the population loss $\mathbb{E}\left\{\displaystyle\lim_{n\to\infty}\mathbb{E}[L^{(s)}_{Cox}(\theta)\lvert D(n)]\right\}$ depend on $s$ and are different from Zhong et al. (2022). Second, when $D(s)$ is generated by randomly sampling $s$ subjects without replacement from $D(n)$ in each iteration, $\mathbb{E}[L^{(s)}_{Cox}(\theta)\lvert D(n)]$ is an average of $\binom{n}{s}$ combinations of mini-batches, where different batches may share the same samples. Hence, the correlation between batches needs to be handled when deriving the properties of the minimizer of $\mathbb{E}[L^{(s)}_{Cox}(\theta)\lvert D(n)]$.
Throughout the paper, we impose the following standard assumptions for the time-to-event data with right censoring:
@@
where $\lVert W_{k}\rVert_{\infty}$ and $\lVert v_{k}\rVert_{\infty}$ denote the sup-norm of a matrix or vector, $\lVert\cdot\rVert_{0}$ is the number of nonzero entries of a matrix or vector, $\lVert f\rVert_{\infty}$ is the sup-norm of the function $f$. The constant $\mathcal{D}>0$ and the sparsity constraint $\varsigma$ is a positive integer. The sparsity assumption is employed due to the widespread application of techniques such as weight pruning (Srinivas et al., 2017), dropout (Srivastava et al., 2014), or $L_{1}$ regularization. These methods effectively reduce the total number of nonzero parameters, preventing neural networks from overfitting, which results in a sparsely connected NN. We approximate $f_{0}$ in (2.1) using a NN $f_{\theta}\in\mathcal{F}(K,\varsigma,\mathbf{p},\mathcal{D})$. More precisely, $f_{0}$ is estimated by minimizing $\mathbb{E}\left[L^{(s)}_{Cox}(\theta)|D(n)\right]$ through the SGD procedure in (2.6), where $D(s)$ is generated by randomly sampling $s$ subjects without replacement from $D(n)$. To simplify the notation, we omit $\theta$ and denote the estimator by
@@
Suppose $\tilde{f}^{(s)}=\displaystyle\arg\min_{f:\lVert f\rVert_{\infty}\leq c,\mathbb{E}[f(X)]=0}L^{(s)}_{0}(f)$ and by definition $L^{(s)}_{0}(\tilde{f}^{(s)})\leq L^{(s)}_{0}(f_{0})$. On the other hand, we have $L^{(s)}_{0}(\tilde{f}^{(s)})\geq L^{(s)}_{0}(f_{0})$ from Lemma 1. Hence, $L^{(s)}_{0}(\tilde{f}^{(s)})-L^{(s)}_{0}(f_{0})=0$ and it implies $\tilde{f}^{(s)}=f_{0}$. That is, for any integer $s\geq 2$, the minimizer of $L^{(s)}_{0}(f)$ on a neighborhood of $f_{0}$ is $f_{0}$ and does not depend on $s$. Our result can be viewed as a generalization of the result in Tarkhan & Simon (2024), where they consider a parametric function $f_{\theta}$ with the truth $f_{0}=f_{\theta_{0}}$ and demonstrate that $\arg\min_{\theta}L^{(s)}_{0}(f_{\theta})=\theta_{0}$.
Next, we establish the consistency of the mb-MPLE $\tilde{f}^{(s)}_{n}$ defined in (3.2) for Cox-NN. Note that the expectation of the estimator $\tilde{f}^{(s)}_{n}$ in (3.2) is not necessarily zero. For the identifiability condition $\mathbb{E}[f_{0}]=0$ in (N2), we propose to apply a mean shift by subtracting the empirical average $\bar{\tilde{f}}^{(s)}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\tilde{f}^{(s)}_{n}(X_{i})$ based on our data. The following theorem gives the convergence rate of the mean-shifted mb-MPLE. This is different from Zhong et al. (2022), where the convergence rate was established for the MPLE after shifting by the true mean.
@@
For any integer $s\geq 2$, the convergence rate of $\tilde{f}^{(s)}_{n}$ is the same as the MPLE in Zhong et al. (2022). The rate is determined by the smoothness and the intrinsic dimension $\mathbf{\tilde{d}}$ of the function $f_{0}$, rather than the dimension $\mathbf{d}$. Therefore, the estimator $\tilde{f}^{(s)}_{n}$ can circumvent the curse of dimensionality (Bauer & Kohler, 2019) and has a fast convergence rate when the intrinsic dimension is low. The batch size $s$ only implicitly influences the constant of the rate. Later, as we consider the parametric Cox model in Section 4, the impact of $s$ on the asymptotic variance of the estimator becomes more apparent.
@@
The minimax lower bound for estimating $f_{0}$ can be derived following the proof of Theorem 3.2 in Zhong et al. (2022) as their Lemma 4 and derivations still hold without the parametric linear term. Specifically, let $\Omega_{0}=\{\lambda_{0}(t):\int_{0}^{\infty}\lambda_{0}(s)ds<\infty\text{ and }\lambda_{0}(t)\geq 0\text{ for }t\geq 0\}$. Under the Cox model with assumptions (A1)-(A3) and (N1), there exists a constant $0<c<\infty$, such that $\inf_{\hat{f}}\sup_{(\lambda_{0},f_{0})\in\Omega_{0}\times\mathcal{H}_{0}}\mathbb{E}\{\hat{f}(X)-f_{0}(X)\}^{2}\geq c\gamma^{2}_{n}$, where the infimum is taken over all possible estimators $\hat{f}$ based on the observed data. Therefore, the NN-estimator in Theorem 1 is rate-optimal since it attains the minimax lower bound up to a polylogarithm factor.

### v3

@@ 1 Introduction
Cox proportional hazards regression (Cox, 1972) is one of the most commonly used approaches in survival analysis, where the outcome of interest is the time to a certain event. It assumes that the covariates have linear effects on the log-hazard function. With the development of deep learning, Cox deep neural networks (Cox-NN, or deep Cox model) have been proposed to capture the potential nonlinear relationship between covariates and survival outcomes to improve survival prediction accuracy (Faraggi and Simon, 1995; Katzman and others, 2018; Ching and others, 2018). Despite the success of the Cox model in survival analysis, it faces a significant optimization challenge when applied to large-scale data. In particular, the Cox model is typically trained by maximizing the partial likelihood (Cox, 1975). The maximum partial likelihood estimator (MPLE) is obtained through the gradient descent (GD) algorithm, which requires the entire dataset to compute the gradient. This approach is computationally demanding and memory-intensive, especially with large datasets. For example, in our motivating application, where images are used to predict survival outcomes (as detailed in Section 6), performing GD in Cox-NN with high-dimensional predictors and a large sample size is infeasible due to hardware memory constraints. Even traditional Cox regression faces optimization challenges for large-scale data. Tarkhan and Simon (2024) reported that the GD algorithm for Cox regression is prone to round-off errors when dealing with large sample sizes. Therefore, the scalability of the Cox model is substantially limited by the inefficiencies of GD optimization when applied to large-scale data.
@@
The stochastic gradient descent algorithm (SGD) is a scalable solution for optimization with large-scale data and has been widely used for training NN (Amari, 1993; Bottou, 2012). SGD alleviates the computational and memory burdens by using a randomly selected (small) subset of the dataset, known as a mini-batch, to compute the gradient and update parameters in each iteration. Moreover, in NN optimization, Xie and others (2020) showed that, compared to GD, SGD favors flat minima, which often leads to better generalization. This explains why SGD is preferred for training an NN. However, SGD cannot directly optimize the partial likelihood of all samples through mini-batches because evaluating the partial likelihood of an event subject requires access to data for all subjects who survived longer.
@@
Several attempts have been made to enable parameter updates for Cox models using mini-batch data. Kvamme et al. (2019) followed the idea of nested case-control Cox regression from Goldstein and Langholz (1992), proposing an approximation of the gradient of partial likelihood using the case-control pairs instead of using all at-risk samples. Sun et al. (2020) fitted a Cox-NN to establish a prediction model for an eye disease progression through SGD, where the iteration was based on the partial likelihood of a random subset of data. Tarkhan and Simon (2024) studied SGD for Cox regression in an online learning setting and demonstrated that the gradient of the expected mini-batch partial likelihood at the true parameter is zero. Despite the successful application of SGD to Cox models, the statistical foundations of the mini-batch maximum partial likelihood estimator (mb-MPLE), of which the SGD seeks to estimate, remain unexplored. Notably, the average mini-batch partial likelihood depends on the batch size and differs from the full partial likelihood, as highlighted in (2.8). Such a distinction in the objective function differentiates mb-MPLE from MPLE, requiring the development of new statistical properties for the mb-MPLE.
@@
  • Secondly, we provide practical guidance for training Cox-NN via SGD to search for the mb-MPLE. In practice, both the learning rate and batch size used in SGD are important hyperparameters, as they significantly influence the SGD dynamics. In optimizations where the objective function is independent of batch size, such as empirical risk minimization, the ratio of the learning rate to the batch size has been identified as a key factor in SGD dynamics during NN training (Goyal and others, 2017; Jastrzebski et al., 2017; Xie and others, 2020). This observation guides a hyperparameter tuning strategy, where the learning rate or the batch size is fixed while tuning the other. However, for Cox-NN, we note that the function targeted by SGD is batch-size dependent, so it remains unclear whether these results can be applied. To this end, we investigate how the local convexity of the objective changes with the batch size and extend the applicability of the tuning strategy to Cox-NN optimization. We provide both theoretical insights and numerical evidence that the ratio of the learning rate to the batch size remains a crucial factor in SGD dynamics during Cox-NN training. This offers insight into Cox-NN hyperparameter tuning.
  • Lastly, we develop the statistical properties of mb-MPLE for Cox regression with linear covariate effects. We show that mb-MPLE is $\sqrt{n}$-consistent and asymptotically normal, with asymptotic variance depending on the batch size. Furthermore, based on theoretical analysis and numerical experiments on the impact of batch size on local convexity, we demonstrate that doubling the batch size improves the statistical efficiency of mb-MPLE. This phenomenon is not observed in other SGD optimizations, such as empirical risk minimization, where the optimizer’s statistical efficiency is independent of the SGD batch size. Moreover, for Cox regression, we study the numerical convergence of the SGD algorithm to the mb-MPLE. We note that the objective function optimized by SGD is not global strong convex, and an additional projection step is necessary (namely, the projected SGD). We follow the non-asymptotic analysis of the projected SGD from Moulines and Bach (2011) to demonstrate that the algorithm approximates the mb-MPLE given sufficiently many iterations.
@@ 2.1 Cox Model and Deep Neural Network
where the matrices $W_{k}\in\mathbb{R}^{p_{k+1}}\times\mathbb{R}^{p_{k}}$ and vectors $v_{k}\in\mathbb{R}^{p_{k+1}}$ (for $k=0,\dots,K$) are the parameters of the NN. The pre-specified activation function $\sigma(\cdot)$ is a nonlinear transformation that operates component-wise on a vector, that is, $\sigma((x_{1},\dots,x_{p_{k}})^{T})=(\sigma(x_{1}),\dots,\sigma(x_{p_{k}}))^{T}$, for $k=1,\dots,K$. For NN in (2.2), $K$ denotes the depth of the network. The sequence $\mathbf{p}$ lists the width of each layer with the first element $p_{0}$ being the dimension of the input variable, $p_{1},\dots,p_{K}$ are the dimensions of the $K$ hidden layers, and $p_{K+1}$ is the dimension of the output layer. The matrix entries $(W_{k})_{i,j}$ are the weight linking the $j$th neuron in layer $k$ to the $i$th neuron in layer $k+1$, and the vector entries $(v_{k})_{i}$ represent a shift term associated with the $i$th neuron in layer $k+1$. In Cox-NN, the output dimension $p_{K+1}=1$ since the output $f_{\theta}(X)$ is a real value. We consider the commonly used activation function named Rectified Linear Units (ReLU) in Nair and Hinton (2010), i.e., $\sigma(x)=\max\{x,0\}$, which has been similarly considered in Schmidt-Hieber (2020) and Zhong et al. (2022).
@@
We refer to the minimizer of (2.3) as the MPLE. In the partial likelihood, sample $i$ who has experienced the event ($\Delta_{i}=1$) is compared to all samples that have not yet experienced the event up to time $T_{i}$, known as the at-risk set. Minimizing (2.3) is typically solved using gradient-descent-based algorithms, where the gradient $\nabla_{\theta}L^{(n)}_{Cox}(\theta)$ is calculated using all $n$ samples at each iteration (Therneau and others, 2015; Katzman and others, 2018; Zhong et al., 2022). The consistency and asymptotic normality of the MPLE have been well-studied by Andersen and Gill (1982) for Cox regression. For Cox-NN, the consistency of the MPLE and its convergence rate have been developed by Zhong et al. (2022). However, when optimizing the Cox model through the SGD algorithm using a subset of samples, the algorithm does not directly minimize (2.3), which will be further explained in the next section.
@@ 2.2 Stochastic Gradient Descent for Cox Model
where $\gamma_{t}$ is a pre-scheduled learning rate for the $t$-th iteration. The gradient ${\nabla_{\theta}}L^{(s)}(D_{t}(s);\hat{\theta}_{t})$ is calculated from the mini-batch data $D_{t}(s)$ at $t$-th iterations. Suppose $D_{t}(s)$ is randomly selected from $D(n)$ with ${\nabla_{\theta}}L^{(s)}(D_{t}(s);\hat{\theta}_{t})$ uniformly bounded for all $t$, when $t\to\infty$, it has been shown that SGD (2.4) with an appropriate learning rate will find the minimizer of $\mathbb{E}_{D_{t}(s)}[L^{(s)}(D_{t}(s);\hat{\theta}_{t})|D(n)]$ if it is strongly convex, where the expectation is over the randomness of generating $D_{t}(s)$ from $D(n)$ (Moulines and Bach, 2011; Toulis and Airoldi, 2017). Throughout the paper, we assume $D_{t}(s)$ is randomly sampled from $D(n)$ at each iteration, so its distribution is independent of $t$. To simplify the notation, for any $\theta$, we omit the input data and drop the subscript $t$ and let $L^{(s)}(\theta):=L^{(s)}(D(s);\theta)$, $L^{(n)}(\theta):=L^{(n)}(D(n);\theta)$, and $\mathbb{E}[L^{(s)}(\theta)|D(n)]:=\mathbb{E}_{D(s)}[L^{(s)}(D(s);\theta)|D(n)]$.
@@
For the Cox model, the SGD implementation (Sun et al., 2020; Tarkhan and Simon, 2024) requires at least two samples to calculate the gradient, which restricts $s\geq 2$. The parameters are updated through
@@ 3.1 Statistical Properties of the mb-MPLE
Zhong et al. (2022) studied the asymptotic properties of the MPLE which minimizes (2.3). They showed that the MPLE achieves the minimax optimal convergence rate up to a polylogarithmic factor. While our work is inspired by the theoretical developments in Zhong et al. (2022), it has significant differences. First, the at-risk term $\frac{1}{n}\sum_{j=1}^{n}I(T_{j}\geq t)\exp\{f_{\theta}(X_{j})\}$ converges to a fixed function $\mathbb{E}[I(Y>t)\exp\{f_{\theta}(X)\}]$. However, in the average mini-batch log-partial likelihood, the at-risk term in $L^{(s)}_{Cox}(\theta)$ always consists of $s$ sub-samples. Thus, both the empirical loss $\mathbb{E}[L^{(s)}_{Cox}(\theta)\lvert D(n)]$ and the population loss $\mathbb{E}[L^{(s)}_{Cox}(\theta)]$ depend on $s$ and are different from Zhong et al. (2022). Second, when $D(s)$ is generated by randomly sampling $s$ subjects without replacement from $D(n)$ in each iteration, $\mathbb{E}[L^{(s)}_{Cox}(\theta)\lvert D(n)]$ is an average of $\binom{n}{s}$ combinations of mini-batches, where different batches may share the same samples. Hence, the correlation between batches needs to be handled when deriving the properties of the minimizer of $\mathbb{E}[L^{(s)}_{Cox}(\theta)\lvert D(n)]$.
We impose the following assumptions for the time-to-event data with right censoring:
@@
where $\lVert W_{k}\rVert_{\infty}$ and $\lVert v_{k}\rVert_{\infty}$ denote the sup-norm of a matrix or vector, $\lVert\cdot\rVert_{0}$ is the number of nonzero entries of a matrix or vector, $\lVert f\rVert_{\infty}$ is the sup-norm of the function $f$. The constant $\mathcal{D}>0$ and the sparsity constraint $\varsigma$ is a positive integer. The sparsity assumption is employed due to the widespread application of techniques such as weight pruning (Srinivas and others, 2017), dropout (Srivastava et al., 2014), or $L_{1}$ regularization. These methods effectively reduce the total number of nonzero parameters, preventing neural networks from overfitting, which results in a sparsely connected NN. We approximate $f_{0}$ in (2.1) using a NN $f_{\theta}\in\mathcal{F}(K,\varsigma,\mathbf{p},\mathcal{D})$. More precisely, $f_{0}$ is estimated by minimizing $\mathbb{E}\left[L^{(s)}_{Cox}(\theta)|D(n)\right]$ through the SGD procedure in (2.6), where $D(s)$ is generated by randomly sampling $s$ subjects without replacement from $D(n)$. To simplify the notation, we omit $\theta$ and denote the estimator by
@@
Suppose $\tilde{f}^{(s)}=\displaystyle\arg\min_{f:\lVert f\rVert_{\infty}\leq c,\mathbb{E}[f(X)]=0}L^{(s)}_{0}(f)$ and by definition $L^{(s)}_{0}(\tilde{f}^{(s)})\leq L^{(s)}_{0}(f_{0})$. On the other hand, we have $L^{(s)}_{0}(\tilde{f}^{(s)})\geq L^{(s)}_{0}(f_{0})$ from Lemma 1. Hence, $L^{(s)}_{0}(\tilde{f}^{(s)})-L^{(s)}_{0}(f_{0})=0$ and it implies $\tilde{f}^{(s)}=f_{0}$. That is, for any integer $s\geq 2$, the minimizer of $L^{(s)}_{0}(f)$ on a neighborhood of $f_{0}$ is $f_{0}$ and does not depend on $s$. Our result can be viewed as a generalization of the result in Tarkhan and Simon (2024), where they consider a parametric function $f_{\theta}$ with the truth $f_{0}=f_{\theta_{0}}$ and demonstrate that $\nabla_{\theta}L^{(s)}_{0}(f_{\theta})|_{\theta=\theta_{0}}=0$.
Next, we establish the consistency of the mb-MPLE $\tilde{f}^{(s)}_{n}$ defined in (3.2) for Cox-NN. Note that the expectation of the estimator $\tilde{f}^{(s)}_{n}$ in (3.2) is not necessarily zero. For the identifiability condition $\mathbb{E}[f_{0}]=0$ in (N2), we apply a mean shift by subtracting the empirical average $\bar{\tilde{f}}^{(s)}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\tilde{f}^{(s)}_{n}(X_{i})$ ba. The following theorem gives the convergence rate of the mean-shifted mb-MPLE. This is different from Zhong et al. (2022), where the convergence rate was established for the MPLE after shifting by the true mean.
@@
For any integer $s\geq 2$, the convergence rate of $\tilde{f}^{(s)}_{n}$ is the same as the MPLE in Zhong et al. (2022). The rate is determined by the smoothness and the intrinsic dimension $\mathbf{\tilde{d}}$ of the function $f_{0}$, rather than the dimension $\mathbf{d}$. Therefore, the estimator $\tilde{f}^{(s)}_{n}$ can circumvent the curse of dimensionality (Bauer and Kohler, 2019) and has a fast convergence rate when the intrinsic dimension is low. The batch size $s$ only implicitly influences the constant of the rate. Later, as we consider the parametric Cox model in Section 4, the impact of $s$ on the asymptotic variance of the estimator becomes more apparent.
@@
The minimax lower bound for estimating $f_{0}$ can be derived following the proof of Theorem 3.2 in Zhong et al. (2022) as their Lemma 4 and derivations still hold without the parametric linear term. Specifically, let $\Omega_{0}=\{\lambda_{0}(t):\int_{0}^{\infty}\lambda_{0}(s)ds<\infty\text{ and }\lambda_{0}(t)\geq 0\text{ for }t\geq 0\}$. Under the Cox model with assumptions (A1)-(A3) and (N1), there exists a constant $c>0$, such that $\inf_{\hat{f}}\sup_{(\lambda_{0},f_{0})\in\Omega_{0}\times\mathcal{H}_{0}}\mathbb{E}\{\hat{f}(X)-f_{0}(X)\}^{2}\geq c\gamma^{2}_{n}$, where the infimum is taken over all possible estimators $\hat{f}$ based on the observed data. Thus, the NN-estimator is rate-optimal as it attains the minimax lower bound up to a polylogarithm factor.